How do you go about solving this problem?
A box with no top is constructed from a rectangular piece of cardboard whose length measures 4 inches more than its width. The box is formed by cutting squares from each corner that measure 2 in. on each side and then folding up the sides. If the volume of the box is 64 cubic in., what are the dimensions of the original piece of cardboard?

Answer by mananth(16946)

Paper
let width of box be x
length of box = x+4
height will be 2 inches since the cut portion is 2 "
width of box = x-4 "
length of box = x+4-4= x
dimensions of new box = x-4, x , 2"
Volume = 2x(x-4)
64=2x^2-8x
2x^2-8x-64=0
/2
x^2-4x-32=0
x^2-8x+4x-32=0
x(x-8)+4(x-8)=0
(x-8)(x+4)=0
x= -4 OR 8
ignore negative
x=width of board 8"
length = 12"